5.A balloon has a volume of 2.25 L when the temperature is 400C. What will be the volume if the balloon is cooled to 120C? Assume the pressure to be constant.


1
Expert's answer
2022-05-11T05:32:04-0400

V1/T1 = V2/T2

2.25 / 400 = V2/120

V2 = 0.68L